OpenStudy (anonymous):

a + 2c - 5b - c + a - b

OpenStudy (anonymous):

help

OpenStudy (igreen):

Combine the like terms, do you know what like terms are?

OpenStudy (anonymous):

no

OpenStudy (igreen):

Like terms are the numbers/letters that are the same. Like in 5 + x + z + 2x x and 2x are like terms, because they both have 'x' as their variable.

OpenStudy (igreen):

Here there are 3 pairs of like terms, can you find them?

OpenStudy (anonymous):

2c and 5b

OpenStudy (anonymous):

Simplify the following expression:

OpenStudy (anonymous):

a + 2c - 5b - c + a - b = a + a - 5b - b + 2c - c = 2a - 6b + c

OpenStudy (igreen):

No. The like terms are: a and a 2c and -c -5b and -b Simplify them by adding/subtracting: a + a = 2a 2c - c = c -5b - b = -6b So we have -6b + 2a + c

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @almightyyzayy12345

OpenStudy (anonymous):

ok thanks

OpenStudy (igreen):

No problem.

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

Ethan is making flat rectangular pizzas for his shop. The number of slices in each pizza is 2 more than the number of pizzas he has on a single countertop. If p represents the number of pizzas on each countertop, which expression can be used to represent the number of slices on 4 countertops?